OpenStudy (anonymous):

need some assistance sin(θ-8π)

OpenStudy (zzr0ck3r):

sin(x) = sin(x+2pi*n)

OpenStudy (zzr0ck3r):

where n is {...-3,-2,-1,0,1,2,3...}

OpenStudy (zzr0ck3r):

so sin(x+8pi) = sin(x+6pi) = sin(x+4pi) = sin(x+2pi) = sin(x)

OpenStudy (zzr0ck3r):

so sin(x-8pi) = sin(x-6pi) = sin(x-4pi) = sin(x-2pi) = sin(x)

OpenStudy (anonymous):

ok but the answer for this question is -2 / square root of 13

OpenStudy (zzr0ck3r):

did they give you a theta?

OpenStudy (zzr0ck3r):

\[sin(\theta+8\pi)=sin(\theta)\]

OpenStudy (zzr0ck3r):

plug in theta to get the answer

OpenStudy (zzr0ck3r):

you with me @darkhound ?

OpenStudy (zzr0ck3r):

sort of a strange answer if the question is sin(theta) = -2/sqrt(13) then theta is a very "strange" angle, of which most people would use a calculator instead of deriving it.

OpenStudy (anonymous):

but the answer is sin(θ-8π)= -2/sqrt(13)

OpenStudy (zzr0ck3r):

write the question exactly as its asked

OpenStudy (zzr0ck3r):

you fall asleep typing?

OpenStudy (anonymous):

given that tanθ=2/3 and θ is in the 3rd quadrant.without evaluating the angle θ,find the exact values of a)cos(-θ) b)sin(-θ) c)sin(θ-8π)

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

lol this is much different than just sin(x-8pi)

OpenStudy (anonymous):

ok but u know how to solve it?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (zzr0ck3r):

soh cah toa says that tan(theta) = opposite over adjacent, and thus you have a triangle like this  so 2^2+3^2=c^2 so c= sqrt(13) so hypotenuse is sqrt(13) soh says sin = opposite of hypotenuse so sin(theta) = -2/sqrt(13) because theta is in quadrant 3

OpenStudy (zzr0ck3r):

hope this helps, brb taco bell:)

OpenStudy (anonymous):

so this is the answer for c right?

OpenStudy (zzr0ck3r):

yes part c) because sin(theta-8pi) = sin(theta)

OpenStudy (zzr0ck3r):

= sin(theta-(2345670*10^53)pi)

OpenStudy (anonymous):



OpenStudy (anonymous):

how come is -2?

OpenStudy (anonymous):

are u still there zzr0ck3r